OpenStudy (anonymous):

FInd the length of curve y=x^3+(1/(12x)) on interval [1/2,2] I get that I take the derivative and square it. and plug it in to sqrt(1+y'^2) but what confuses me is do I do the inverse of derivative to equal out. also can I just take the square root of everything inside and then take the integral?

OpenStudy (freckles):

you can take the square root of the thing inside if the inside itself can be written as a square but depending on the limits you might have to adjust your sign of the integrand

OpenStudy (freckles):

so for this one what do you get as y'?

OpenStudy (anonymous):

I get \[\sqrt{1+((3x^2-(1/12x^2))^2}\] so could I basically just cancel out the square root

OpenStudy (freckles):

no no

OpenStudy (freckles):

the whole thing inside has to be written as a square for you to do that

OpenStudy (freckles):

like so: \[1+(3x^2-\frac{1}{12x^2})^2 \\ 1+(9x^4-\frac{1}{2}+\frac{1}{144x^2}) \\ 9x^4+\frac{1}{2}+\frac{1}{144x^2} \\ (3x^2+\frac{1}{12x^2})^2 \]

OpenStudy (freckles):

\[\sqrt{(3x^2+\frac{1}{12x^2})^2}=3x^2+\frac{1}{12x^2} \text{ for all } x \neq 0\]

OpenStudy (freckles):

which we don't have to worry about 0 since we are looking at [1/2,2]

OpenStudy (freckles):

so this is what you have: \[\int\limits_{1/2}^{2} (3x^2+\frac{1}{12x^2}) dx\]

OpenStudy (freckles):

made a type-o

OpenStudy (freckles):

\[1+(3x^2-\frac{1}{12x^2})^2 \\ 1+(9x^4-\frac{1}{2}+\frac{1}{144x\color{red}{^4}}) \\ 9x^4+\frac{1}{2}+\frac{1}{144x\color{red}{^4}} \\ (3x^2+\frac{1}{12x^2})^2 \]

OpenStudy (freckles):

by the way the whole sign thing I mentioned earlier...pretend we have: \[\int\limits_{-1}^2 \sqrt{x^2} dx \\ \text{ well } \sqrt{x^2}=x \text{ if } x \ge 0 \\ \text{ and } \sqrt{x^2}=-x \text{ if } x<0 \\ \text{ so } \\ \int\limits_{-1}^2 \sqrt{x^2} dx=\int\limits_{-1}^0 (-x )dx+\int\limits_0^2 x dx\]

OpenStudy (anonymous):

So in reality there is no square root for this problem? because the two integrals are equal to it?

OpenStudy (freckles):

yeah we got rid of the square root by writing previous integrand as \[3x^2+\frac{1}{12x^2}\]

OpenStudy (anonymous):

didn't you forget the plus 1?

OpenStudy (freckles):

remember this: \[1+(3x^2-\frac{1}{12x^2})^2 \\ 1+(9x^4-\frac{1}{2}+\frac{1}{144x^4}) \\ 9x^4+\frac{1}{2}+\frac{1}{144x^4} \\ (3x^2+\frac{1}{12x^2})^2\]

OpenStudy (freckles):

those 4 expressions are equal expressions

OpenStudy (freckles):

\[\sqrt{1+(3x^2-\frac{1}{12x^2})^2} =\\ \sqrt{(3x^2+\frac{1}{12x^2})^2}\]

OpenStudy (freckles):

hey @Empty and @IrishBoy123 I have to leave can you help @Greenwalrus further

OpenStudy (anonymous):

ah yeah I see!

OpenStudy (anonymous):

help :(

OpenStudy (anonymous):

since I don't have a negative I can just ignore having two integrals and have one right